OpenStudy (anonymous):

Students in an English class took a final exam. They took equivalent forms of the exam at monthly intervals thereafter. The average score S(t), in percent, after t months was found to be given by S(t) = 68 − 20 log (t + 1), t ≥ 0. What was the average score after 14 months?

OpenStudy (anonymous):

t = 14. Sub in t = 14 into S(t) = 68 − 20 log (t + 1), S(t) = 68 − 20 log (14 + 1) = S(t) = 68 − 20 log (15) ...